xxWorkshop on Open Source Computer Algebra

Workshop on Open Source Computer Algebra

Tuesday 21st - Thursday 23rd of May 2002 Lyon, France

Amphi G3, Bxtiment enseignement [A] University Claude-Bernard LyonxI, Campus de Gerland 50, avenue Tony-Garnier, Domaine de Gerland, 69366 Lyon Cedex 07

Organizing committee: Daniel Duparc, Bernard Mourrain, Bernard Parisse, Fabrice Rouillier, Marie-Franxoise Roy, Nicolas Thixry, and Paul Zimmermann.

Local organizers: Nicolas Thixry, Gilles Villard Contact: nthiery@users.sf.net

Web page: http://www.lapcs.univ-lyon1.fr/~nthiery/CalculFormelLibre/workshop.html

PIC PIC

This workshop was organized with the support of the GDR Algorithme, Langage et Programmation, of the AS "Calcul Formel", of the Laboratoire de l’Informatique du Parallxlisme (ENS-Lyon), of the INRIA, and of the Union des Professeurs de Spxciales.

Here is a preliminary version of the report on this workshop.

1 Scope The organization of this workshop was initiated after long discussions on the French mailing list calcul-formel-libre@math.cnrs.fr last autumn, where the lack of an open source computer algebra system was deplored. See

http://www.math.cnrs.fr/archives/logiciels-formels/ for the archives, and
http://www.lapcs.univ-lyon1.fr/~nthiery/CalculFormelLibre/ for a tentative

summary (in French).

The aim of this workshop is to gather potential users and developers from academy and industry, in order to:

Take a snapshot of the current situation of free software for computer algebra;

Specify the needs, for research, teaching, and industry usage as well;

Study the feasibility of a free system in this domain.

Here, by free we mean open source, and more precisely compatible with the GPL (see

http://www.fsf.org/licenses/license-list.html#GPLCompatibleLicenses); by

system we mean indifferently a software or a collection of software.

2 Registration To register to the workshop, please send an e-mail to Helene Zganic (Helene.Zganic@loria.fr), with the following information:

Name, title:

Position:

Address:

E-mail:

Tel, fax:

Expected arrival and departure dates:

Attendance to the banquet [yes/no]:

If you plan to attend the banquet, please send your registration form before may

10th!

We basically need to know the number of participants in order to organize the coffee breaks, the lunches, and the banquet. There is no registration fee. The banquet will be offered, or proposed at a low price if there are more than 50 participants. The lunches will be taken at the ENS-Lyon. The participants have to take care for themselves for the lodging. You can find some suggestions on

http://www.loria.fr/projets/spaces/cfl.html.

3 Tentative program Poster sessions will take place during the whole workshop. If you would like to present a poster about some software, please send us a description before Mai 15th.

Tuesday May 21st, afternoon Theme: definition of the goals and constraints

Opening session

Open source software, licenses, and legal issues: Bernard Lang

Coffee break

TeXmacs Joris van der Hoeven

Software demonstrations:

GINAC: Richard B. Kreckel Slides GIAC: Bernard Parisse Slides Linbox Pascal Giorgi & Gilles Villard TRIP: Jacques Laskar et Mickael Gastineau ACE, mu-EC, and mupad-combinat: Florent Hivert

All the participants are welcome to join the organizers for dinner at the nearby Ninkasi restaurant, 267 r Marcel Mxrieux 69007 LYON.

Wednesday May 22nd, morning Theme: Development Tools and Models

Objective Caml: Pierre Weis SYNAPS: P. Trxbuchet Coffee break Scilab: Claude Gomez Slides GAP System development for Computer Algebra: Alexander Hulpke Meal at the nearby xcole Normale Supxrieure de Lyon

Wednesday May 22nd, afternoon Theme: Existing platforms and software MuPAD 2.5 - News, Trends and Outlooks: Oliver Kluge

The MAGMA Computer Algebra System: Allan Steel, University of Sydney

Axiom Tim Daly

Coffee break

Maxima: Annick Valibouze; presented by Marc Giusti

Foc: Renaud Rioboo

PARI: Karim Belabas Slides

Singular: Pfister Slides

All the participants are invited for the banquet at the Brasserie Georges, 30, cours de Verdun Perrache, just nearby gare Perrache.

Thursday May 23rd

g++: Gabriel Dos Reis

This final day will gather the participants who want to get personally involved in a "free computer algebra system" project. Volunteers are warmly welcome! Please note that this does not necessarily require special technical skills, as much of the work will consist in beta-testing and writing documentation. The goal is to write down a synthesis document for this workshop, to define the main lines of the project, and last but not least to find volunteers to contribute!

Session 1: Potential users and their needs. students, teachers and professors researchers (academic) engineers (industry)

Coffee break Session 2: Technical points.

specify some precise components and their interfaces, for example: kernel, library, common platform/environment, compiler, documentation, test suite, benchmarks, ... for each component: can we restart from existing free software, or push an existing software to become free? for each interface: can we reuse an existing (standard or popular) language? connection with existing (non necessarily free) systems? how to include contributions: referee process by publishing board? Define rules for contributions: benchmarks, bug corrections, ... diffusion of the whole distribution: cvs server, ftp, cdrom? which language for documentation: pdf?

(In parallel with session 2): Working groups. Interface with numerical libraries (NAG, Octave, Scilab, GnuGSL) [moderator Tim Daly] Corba-like communication protocol/server for computer algebra systems [moderator Joris van der Hoeven] Meal at the nearby xcole Normale Supxrieure de Lyon Session 3: Project management.

legal aspects: which law applies (european/US)? financial aspects manpower, how to find it? who decides what: executive committee? strategic choices (license, form of collaborative development) ensure the diffusion of the system: commercial partners?

Coffee break Session 4: Workplan.

list some well-defined tasks and people to achieve them create one mailing-list? a web page? evaluate the required funding and how/where to find it fix deadlines for each technical project? constitute an executive committee and a publishing board

End of the workshop

4 Guidelines for presenting a software Here are some questions that we would like you to answer about the software you are going to present. Please be helpful: avoid hype, and be as neutral as you can!

What are the main strengths and weaknesses of the software ? What development model is used ? What difficulties are encountered ? What are the reasons for the success or failure of the software ? What’s the future of the software ? How is the software distributed ? What kind of license is used for the software ? Why ? Is this an open-source license? If not, would it be possible to switch to an open-source license ? Is there an active community of users ? Is there an active community of contributers ? How are contributions integrated and rewarded ? Are there contributions for the core of the system, or only as external modules ? What would happen if you stopped the development of the software ? Is there documentation (for developers / for users) ? Who writes it ? Are there books with tutorials, exercises, and examples ? Are they written by developers of the project or external persons ? Where is the software used ? university, industry, high school/university teaching, research, ...

What is the size of the developer team ?

5 How to come ? Coming by car: well, just check a map of Lyon; you cannot miss avenue Tony Garnier on it (south of Lyon, on the east side of the Rhxne). But prepare for some trouble finding a parking lot.

Coming by plane: there is an international airport in Lyon Satolas, with a bus shuttle that takes you downtown to the railway station "Lyon Part-Dieu" (around 20 euros round trip). Alternatively, there are direct fast trains (TGV, 2 hours for 500 kms) on a very regular basis between the international airport Paris Charles-de-Gaule and Lyon. The price of the train is about 120 euros round trip.

Coming by train: there are two main railway stations in Lyon, many trains going through both of them. We suggest that you use "Lyon Part-Dieu", though "Lyon Perrache" is not much further away from the campus. See http://voyages.sncf.fr/ for schedules, fares and reservations.

Now, coming from the railway station "Lyon Part-Dieu" to the campus is rather easy. Just beware that this is not the main campus "La Doua" of the university Lyon I. The workshop indeed takes place in new buildings, nearby the xcole Normale Supxrieure de Lyon.

Take the subway, line B, all the way down to "Stade de Gerland" (south of Lyon). At the exit, walk to the west, on avenue Tony Garnier, about 200 meters past the gas station; there are constructions so only the pavement on the right of the avenue is walkable. Cross the street at the light. The building [A] is the 3 stories building with a wood and glass structure that stands in front of you.

You can also check the web site of the Transports Communs de Lyon http://www.tcl.fr/.